OpenStudy (anonymous):

Plz help...I'll be ur fan................ Blake goes running for 10 minutes on Sunday. Every day he doubles his running time. If he continues this pattern, how many minutes will he run on the 7th day? I'm confused about which formula I should use: f(n) = f(1) • r^(n−1) or an = a1 • r^(n−1) are these even the correct formulas for this question????

OpenStudy (anonymous):

@moazzam07

OpenStudy (rajat97):

yes those are the correct ones and both say the same thing if you use the second one, a1=10, r=2 and n=7

OpenStudy (anonymous):

So that means that on the 7 day he would run 640 minutes???? @rajat97

OpenStudy (rajat97):

i'm calculating!

OpenStudy (rajat97):

yes you are right

OpenStudy (princeharryyy):

what do you mean by doubling

OpenStudy (rajat97):

no that's not right @princeharryyy has got it

OpenStudy (princeharryyy):

day 1 =10 =10 day2 = 2*10 =20 day3 = 2*2*10 =40 ..... day7 = (2^6 )* 10 = 64*10 =640minutes @rajat97 and @plzzhelpme

OpenStudy (princeharryyy):

and that was right @rajat97

OpenStudy (anonymous):

wait........it was right..

OpenStudy (rajat97):

oh yeah! silly me!

OpenStudy (princeharryyy):

yeah! silly you! that was right.

OpenStudy (anonymous):

ok thanks alot... what formula would u use for this question?? Alex studies for an exam for 50 minutes on Monday. Every day he increases his study time for 5 minutes. If he continues this pattern, how many minutes will he study on the 7th day?

OpenStudy (rajat97):

this is an arithematic progression

OpenStudy (rajat97):

you can use a1 + (n-1)d

OpenStudy (rajat97):

here d=5, n=7and a1=50

OpenStudy (princeharryyy):

50, 50+5, 50+5+5, 50+5+5+5, ..... and so on

OpenStudy (anonymous):

and that would be 80 minutes on day 7 right?

OpenStudy (rajat97):

yup

OpenStudy (anonymous):

ok....thanks alot:* :* :*

OpenStudy (princeharryyy):

* :* :*